The senior class at high school A rented and filled 1 van and 6 buses with 372 students. High school B rented and filled 4 vans and 12 buses with 780 students. Each van and each bus carried the same number of students. How many students can a van carry? How many students can a bus carry?
Found 2 solutions by mananth, sudhanshu_kmr:
Answer by mananth(16949)

let van carry x students
bus carry y students
School A = x+6y=372...............1
School B = 4x+12y=780.............2
..
multiply equation 1 by -2
-2x-12y=-744
add this equation to 2
2x=36
x= 18 capacity of van
plug value of x in equation 1
18+6y=372
6y=372-18
6y=354
y=59 capacity of bus
...
m.ananth@hotmail.ca

Answer by sudhanshu_kmr(1152)

Let each van carry x students, and
each bus carry y students.

according to question.
x + 6y = 372........(1)
4x +12y = 780 .....(2)
now, multiply (1) by 4 subtract from it (2)

4x + 24 y = 1488
4x + 12 y = 780
then 12 y = 708 , y =708/12 = 59
and x = 372 - 354 = 18


a van will carry 18 students
a bus will carry 59 students......




It is possible that some typing mistake in solution of a problem, if any please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......